# Impact of LLMs news Sentiment Analysis on Stock Price Movement Prediction

### v1

@@ 2 Proposed Experimental Study
The study examines five major publicly traded companies – Microsoft (MSFT), Amazon (AMZN), Apple (AAPL), Netflix (NFLX), and Tesla (TSLA) – over the period from March $10^{th}$ 2022 to April $2^{nd}$ 2025. We collect the related daily stock data from Yahoo Finance and the corresponding news (total of over 96000) via the AlphaVantage API. News are assigned to trading days according to their release timestamps, with weekend and holiday publications mapped to the next available trading day. Any missing values in the price series are handled through forward filling for continuity in time-series modeling.
In order to evaluate the efficiency, complementarity of the Financial sentiment analysis models and to train the ensemble models, we have used a labelled financial news dataset; SEntFiN 1.0 [10]. Unlike the Alphavantage dataset, SEntFiN provides news as text-sentiment pairs (over 10,700 news headlines), enabling model evaluation, but without any information about their issuing date, or related stock, which makes it unsuitable for the rest of the study.
@@
are defined in four forms: price, factor, delta, and binary trend movement. Formally, let $P_{t}$ denote the closing price at trading day $t$, and let $\tilde{P}_{t}=P_{t+1}$ represent the next-day closing price aligned to time $t$. The factor (price ratio) is computed as $\text{Factor}_{t}=\frac{\tilde{P}_{t}}{P_{t}}=1+r_{t}.$, the delta (price change) as $\Delta_{t}=\tilde{P}_{t}-P_{t}=P_{t+1}-P_{t}.$ and the Binary (upward vs. downward movement) as $\text{Binary}_{t}=\mathbb{1}[\tilde{P}_{t}>P_{t}]=\begin{cases}1,&\tilde{P}_{t}>P_{t},\\ 0,&\tilde{P}_{t}\leq P_{t}.\end{cases}$
@@
Predictive performance is quantified using classification and regression metrics, depending on the target formulation. For instance, to evaluate the efficiency in predicting directional (up or down) stock trends, we compute the Accuracy, F1-score, and the Area Under the ROC Curve (AUC). For regression tasks, i.e. predicting the stock price change, we compute the Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and the coefficient of determination $R^{2}$ to approximate the variance proportion explained by the model.
@@
We conduct the experiments on a standard computing environment [^1] NVIDIA GeForce GTX 1650 Ti GPU with 4 GB VRAM, a 2.60 GHz CPU and 16 GB of RAM. The system is equipped with CUDA 12.7 and driver version 565.77.01.. We implement the experimental study pipelines in Python using well-known open-source data manipulation and machine learning libraries. We make our complete codebase and setup instructions with additional details available  [^2] https://github.com/Walids35/capstone-stock-prediction to enable reproducibility and support similar future research.
@@ 3.1 Effectiveness and complementarity of sentiment analysis models
Although being a generic-purpose LLM, DeBERTa outperforms both RoBERTa and FinBERT, predicting sentiments accurately for over 75% of the dataset, followed by FinBERT then RoBERTa scoring both below 70%; 70.9% and 58.9%, respectively. The overlap and disjoint sets of accurately predicted news presented in the Venn diagrams of Figure 1 endorse the complementarity between the models, showing that they do not accurately predict the same news of the same classes. This motivates the training of a learning approach which would learn how to combine results from each model in order to improve the sentiment prediction performance. The results reported in Table 1 endorse this idea, showing an advantage in employing an ensemble model over using any LLM separately, i.e. an SVM ensemble model achieves on average about 80% of accuracy, precision, recall and F1 score.
@@
3.2 Impact of news sentiment on stock movement prediction
[…]
LSTM WSavg 0.5747 / 0.5557 6.1331 / 1.4726 0.0334 / 2.1242 15.8240 / 0.7455
LSTM NS 0.5694 / 0.5453 6.0629 / 1.4615 0.0324 / 2.0703 12.0457 / 0.5769
LSTM WSsvm 0.5720 / 0.5530 6.0971 / 1.4642 0.0330 / 2.1033 14.2921 / 0.6817
PatchTST WSavg 0.5525 / 0.4997 0.2638 / 2.9300 0.2227 / 3.1490 0.3768 / 3.6745
PatchTST NS 0.5558 / 0.5105 0.3474 / 3.5419 0.3906 / 5.0298 0.2419 / 2.3809
PatchTST WSsvm 0.5593 / 0.5137 0.3056 / 3.2822 0.2052 / 2.8807 0.3400 / 3.3225
TimesNet WSavg 0.5656 / 0.5191 0.2609 / 2.8192 0.2934 / 3.9363 0.2932 / 3.0241
TimesNet NS 0.5683 / 0.5209 0.4398 / 4.5547 0.4733 / 6.2774 0.1767 / 1.9371
TimesNet WSsvm 0.5657 / 0.5099 0.2663 / 2.7926 0.3063 / 4.0162 0.2648 / 2.7362
tPatchGNN WSavg 0.5242 / 0.4914 0.1423 / 1.5588 0.1771 / 2.2583 0.2734 / 2.7610
tPatchGNN NS 0.5165 / 0.4664 0.1367 / 1.5147 0.1703 / 2.1881 0.1971 / 2.0581
tPatchGNN WSsvm 0.5117 / 0.4866 0.1426 / 1.5617 0.1782 / 2.2708 0.2559 / 2.5903
In Table 2 we report the average stock movement prediction results of the studied architectures, achieved without news sentiment features (NS), using the SVM ensemble model predicted sentiments (WSsvm) and the average performance when using sentiment features of the 6 considered sentiment analysis classifiers (WSavg). Overall, the results show that the sentiment news has very low impact on the obtained results. Moreover, we notice that some architectures are better fitted for classification (up or down trend prediction), and others for regression (price prediction). For instance, LSTM predicts the movement trends the most accurately (57.47%), while it performs the worst in the regression tasks, yielding the highest MAEs. Overall, we observe that the sentiment features improve (slightly) the stock movement predictions of models trained on the studied architectures, i.e. PatchTST and TimesNet in regression tasks and all classifiers except TimesNet-based ones.

### v2

@@ 2 Proposed Experimental Study
The study examines five major publicly traded companies – Microsoft (MSFT), Amazon (AMZN), Apple (AAPL), Netflix (NFLX), and Tesla (TSLA) – over the period from March $10^{th}$ 2022 to April $2^{nd}$ 2025. We collect the related daily stock data from Yahoo Finance and the corresponding news (total of over 96000) via the AlphaVantage API. News are assigned to trading days according to their release timestamps, with weekend and holiday publications mapped to the next available trading day. Any missing values in the price series are handled through forward filling for continuity in time-series modeling. In order to evaluate the efficiency, complementarity of the Financial sentiment analysis models and to train the ensemble models, we have used a labelled financial news dataset; SEntFiN 1.0 [10]. Unlike the Alphavantage dataset, SEntFiN provides news as text-sentiment pairs (over 10,700 news headlines), enabling model evaluation, but without any information about their issuing date, or related stock, which makes it unsuitable for the rest of the study.
@@
are defined in two forms: price factor (regression task) and binary trend movement (classification task). Formally, let $P_{t}$ denote the closing price at trading day $t$, and let $\tilde{P}_{t}=P_{t+1}$ represent the next-day closing price aligned to time $t$. The factor (price ratio) is computed as $\text{Factor}_{t}=\frac{\tilde{P}_{t}}{P_{t}}=1+r_{t}.$ and the Binary (upward vs. downward movement) as $\text{Binary}_{t}=\mathbb{1}[\tilde{P}_{t}>P_{t}]=\begin{cases}1,&\tilde{P}_{t}>P_{t},\\ 0,&\tilde{P}_{t}\leq P_{t}.\end{cases}$
@@
Predictive performance is quantified using classification and regression metrics, depending on the target formulation. For instance, to evaluate the efficiency in predicting directional (up or down) stock trends, we compute the F1-score, and the Area Under the ROC Curve (AUC). For regression tasks, i.e. predicting the stock price change, we compute the Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE)to approximate the variance proportion explained by the model.
@@
We conduct the experiments on a standard computing environment [^1] NVIDIA GeForce GTX 1650 Ti GPU with 4 GB VRAM, a 2.60 GHz CPU and 16 GB of RAM. The system is equipped with CUDA 12.7 and driver version 565.77.01.. We implement the experimental study pipelines in Python using well-known open-source data manipulation and machine learning libraries. We make our complete codebase and setup instructions with additional details available  [^2] https://github.com/aiubdiuabzdiuba/ICLR2026 to enable reproducibility and support similar future research.
@@ 3.1 Effectiveness and complementarity of sentiment analysis models
Although being a generic-purpose LLM, DeBERTa outperforms both RoBERTa and FinBERT, predicting sentiments accurately for over 75% of the dataset, followed by FinBERT then RoBERTa scoring both below 70%; 70.9% and 58.9%, respectively.
@@
The overlap and disjoint sets of accurately predicted news presented in the Venn diagrams of Figure 1 endorse the complementarity between the models, showing that they do not accurately predict the same news of the same classes. This motivates the training of a learning approach which would learn how to combine results from each model in order to improve the sentiment prediction performance. The results reported in Table 1 endorse this idea, showing an advantage in employing an ensemble model over using any LLM separately, i.e. an SVM ensemble model achieves on average about 80% of accuracy, precision, recall and F1 score.
3.2 Comparison of different daily news sentiment aggregation methods
To assess the contribution of sentiment representations (aggregation of daily news), we analyze the performance of LSTM when different sentiment features are incorporated into market data. Specifically, we design four ablation experiments to isolate the effect of each sentiment type: LSTM (using all sentiment features), LSTM_wo_count (excluding news count per label), LSTM_wo_sum (excluding sentiment score aggregation), LSTM_wo_count_sum (excluding both news count and sentiment aggregation), and LSTM_wo_majority (excluding majority vote aggregation). The performance of these variants is then evaluated across all output types.
For the binary price prediction task (Table  2, the baseline LSTM model achieves the best overall balance across metrics, with an AUC of 0.5557, Accuracy of 0.5747, F1-Score of 0.5507, Precision of 0.6042, and Recall of 0.5747. The ablation variants generally result in reduced performance, confirming the importance of each component. Specifically, removing the count feature (LSTM_wo_count) lowers both AUC (0.5393) and F1-Score (0.5361), while the sum feature also contributes positively, as its removal (LSTM_wo_sum) decreases AUC to 0.5468 and F1-Score to 0.5411. The combination removal (LSTM_wo_count_sum) produces the weakest outcome with an AUC of 0.5318 and F1-Score of 0.5251, showing the compounding effect of feature reduction. Interestingly, excluding the majority component has minimal impact, with LSTM_wo_majority maintaining similar performance to the baseline. These results highlight that while the baseline LSTM provides the strongest predictive power, count and sum features play a key role in supporting accurate classification.
Variant AUC F1-Score
LSTM 0.5557 0.5507
LSTM_wo_count 0.5393 0.5361
LSTM_wo_count_sum 0.5318 0.5251
LSTM_wo_sum 0.5468 0.5411
LSTM_wo_majority 0.5473 0.5495
Table 2: Binary price prediction performance for LSTM and its ablation variants.
3.3 Effectiveness of sentiment analysis models on stock movement classification
Table 3: Performance comparison of sentiment analysis models on stock movement trend prediction (binary classification: stock price going up or down). NS stands for ”No Sentiment” and refers to a baseline stock prediction without any news sentiment feature.
LSTM PatchTST TimesNET tPatchGNN
Model F1 (avg, sd) AUC (avg, sd) F1 (avg, sd) AUC (avg, sd) F1 (avg, sd) AUC (avg, sd) F1 (avg, sd) AUC (avg, sd)
NS 0.541 ± 0.078 0.545 ± 0.066 0.519 ± 0.074 0.511 ± 0.048 0.534 ± 0.087 0.521 ± 0.077 0.430 ± 0.114 0.466 ± 0.060
FinBERT 0.564 ± 0.057 0.562 ± 0.055 0.507 ± 0.074 0.487 ± 0.052 0.544 ± 0.065 0.529 ± 0.053 0.431 ± 0.112 0.491 ± 0.049
DeBERTa 0.534 ± 0.075 0.547 ± 0.038 0.485 ± 0.079 0.491 ± 0.057 0.540 ± 0.068 0.533 ± 0.057 0.444 ± 0.108 0.498 ± 0.052
RoBERTa 0.561 ± 0.039 0.564 ± 0.037 0.502 ± 0.087 0.495 ± 0.056 0.538 ± 0.064 0.514 ± 0.053 0.443 ± 0.108 0.478 ± 0.061
LR 0.552 ± 0.062 0.559 ± 0.046 0.511 ± 0.081 0.507 ± 0.057 0.536 ± 0.057 0.522 ± 0.054 0.461 ± 0.113 0.500 ± 0.063
RF 0.539 ± 0.054 0.549 ± 0.037 0.508 ± 0.070 0.505 ± 0.054 0.509 ± 0.074 0.506 ± 0.049 0.449 ± 0.114 0.495 ± 0.062
SVM 0.554 ± 0.049 0.553 ± 0.041 0.519 ± 0.076 0.514 ± 0.057 0.528 ± 0.061 0.510 ± 0.050 0.411 ± 0.109 0.487 ± 0.054
Table 3 reports the performance of the different stock prediction models (columns) when fed with daily news predicted-sentiment features (rows), in terms of average F1-score, AUC and their corresponding standard deviation computed accross 10 different runs.
The results depict also that LSTM and TimesNET obtain higher performance values compared to PatchTST and tPatchGNN across most configurations. Under the LSTM architecture, the highest F1-score is achieved by SVM (0.554 ± 0.049), while the highest AUC is obtained by FinBERT (0.562 ± 0.055). For PatchTST, SVM records the best performance in both F1 (0.519 ± 0.076) and AUC (0.514 ± 0.057). Within TimesNET, DeBERTa achieves the highest F1 (0.540 ± 0.068) and AUC (0.533 ± 0.057). In contrast, tPatchGNN shows lower values overall, with Logistic Regression obtaining the highest F1 (0.461 ± 0.113) and AUC (0.500 ± 0.063) within that architecture.
When comparing sentiment-based models (FinBERT, DeBERTa, RoBERTa) to the no-sentiment (NS) baseline, increases in AUC are observed for LSTM with FinBERT (0.562 vs. 0.521) and for TimesNET with DeBERTa (0.533 vs. 0.521). In other cases, the differences between sentiment-based models and the NS baseline remain relatively small across architectures.
This indicates, that although SVM stacking approach predicts more accurately sentiments in the FPB dataset (Table 1), it does not always provide the best guidance fo stock movement prediction for every approach. Moreover, models do not benefit equally from the added news sentiment features. Interestingly, we notice that LSTM and PatchTST are more suited for the classification task than TimesNET and tPatchGNN.
3.4 Effectiveness of sentiment analysis models on stock movement regression
Table 4: Performance comparison of sentiment analysis models on stock price movement prediction; regression task on the price factor. NS stands for ”No Sentiment” and refers to a baseline stock prediction without any news sentiment feature.
LSTM PatchTST TimesNET tPatchGNN
Model MAE (avg, sd) RSE (avg, sd) MAE (avg, sd) RSE (avg, sd) MAE (avg, sd) RSE (avg, sd) MAE (avg, sd) RSE (avg, sd)
NS 0.032 ± 0.009 2.070 ± 0.663 0.391 ± 0.186 5.030 ± 2.138 0.483 ± 0.215 6.399 ± 2.593 0.170 ± 0.085 2.188 ± 0.750
FinBERT 0.033 ± 0.010 2.116 ± 0.694 0.251 ± 0.166 3.493 ± 2.293 0.274 ± 0.145 3.769 ± 1.794 0.176 ± 0.088 2.247 ± 0.781
RoBERTa 0.034 ± 0.011 2.180 ± 0.727 0.237 ± 0.149 3.384 ± 2.064 0.254 ± 0.158 3.620 ± 2.197 0.176 ± 0.087 2.244 ± 0.779
DeBERTa 0.033 ± 0.010 2.098 ± 0.671 0.208 ± 0.148 3.020 ± 2.113 0.291 ± 0.146 3.967 ± 1.840 0.175 ± 0.087 2.239 ± 0.777
LR 0.033 ± 0.010 2.112 ± 0.689 0.218 ± 0.133 3.038 ± 1.928 0.298 ± 0.157 3.914 ± 1.926 0.178 ± 0.088 2.270 ± 0.791
RF 0.034 ± 0.010 2.136 ± 0.705 0.217 ± 0.125 3.078 ± 2.152 0.337 ± 0.179 4.332 ± 2.208 0.179 ± 0.088 2.279 ± 0.794
SVM 0.033 ± 0.010 2.103 ± 0.678 0.205 ± 0.122 2.881 ± 1.580 0.306 ± 0.165 4.016 ± 2.076 0.178 ± 0.088 2.271 ± 0.791
Table 4 reports the performance of the different stock prediction models (columns) when fed with daily news predicted-sentiment features (rows), in terms of average MAE, RSE and their corresponding standard deviation computed accross 10 different runs. Lower error values imply better performance.
The influence of sentiment features on forecasting accuracy varies noriceably across architectures. LSTM and tPatchGNN exhibit minimal sensitivity to sentiment integration, with differences between the NS baseline and sentiment-augmented models remaining within 0.002-0.009 for MAE and 0.051-0.110 for RSE, and both architectures achieving their lowest errors without sentiment features. In contrast, PatchTST and TimesNET demonstrate substantial performance gains from sentiment incorporation. For PatchTST, sentiment models reduce MAE by 0.140-0.186 and RSE by 1.537-2.149 compared to the NS baseline (0.391 MAE, 5.030 RSE), with SVM-based sentiment yielding the largest improvement (0.205 MAE, 2.881 RSE).
TimesNET shows even more pronounced benefits, where sentiment integration reduces MAE by 0.146-0.229 and RSE by 2.067-2.779 relative to NS (0.483 MAE, 6.399 RSE), with RoBERTa producing the strongest gains (0.254 MAE, 3.620 RSE). Across sentiment modeling approaches, traditional ML methods (SVM, LR) and transformer-based models (DeBERTa, RoBERTa, FinBERT) show comparable effectiveness within each architecture, with no single sentiment extraction method consistently outperforming others across all forecasting models.